# Emergence of Meron Kekulé lattices in twisted Néel antiferromagnets

### v1

@@ Results
When replicated across multiple supercells to establish an energy-minimized configuration, Meron hexads form a regular triangular lattice that overlaps with the underlying moiré superlattice. However, the constituent meron cores of the Meron hexads create a more intricate pattern, where two nearest-neighbor meron cores from the same Meron hexad have a shorter distance than those two from adjacent Meron hexads (Figure 3a). This distorted structure is often referred to as a Kekulé lattice with a Kekulé-O pattern in the literature 28. Thus, we term this distorted meron lattice a “Meron Kekulé lattice (MK).” The Kekulé structure of the MK state is characterized by two key length parameters: the intracell bond length within the same Meron hexad ($l$) and the intercell bond length across different Meron hexads ($l^{\prime}=L-2l$), with $L$ representing the lattice constant of the moiré superlattice (Figure 4a). The distinction between $l$ and $l^{\prime}$ indicates the deviation from the conventional honeycomb lattice, giving rise to an alternative Kekulé structure, as schematically illustrated in Figure 4a. To quantify this deviation, we introduce a dimensionless constant $\lambda$, referred to as the “Kekulé constant:”
$\lambda=\frac{l^{\prime}-l}{L}=1-\frac{3l}{L}.$ (7)
If $\lambda=0$, the meron lattice aligns with a regular honeycomb lattice. Conversely, if $\lambda\neq 0$, it corresponds to a Kekulé lattice. Typically, $l$ is less than $l^{\prime}$, leading to $\lambda>0$ (e.g., $\lambda=0.19$ for the meron lattice configuration in Figure 3).
Three length parameters ($l$, $l^{\prime}$, and $L$) relevant to the MK state are primarily influenced by specific characteristics of the AFM patch, including its spatial size ($r$) and interlayer exchange coupling strength within the patch. The minimization of interlayer exchange energy favors the alignment between $l$ and $r$, as having $l\approx r$ facilitates an energy-minimized AFM domain array configuration. Conversely, other aspects in the spin model, such as intralayer exchange interactions, interact with the interlayer exchange coupling of the AFM patch, potentially causing deviations from $l=r$. Thus, there exists an intriguing possibility that by manipulating various magnetic interactions in twisted antiferromagnets—represented by parameters like $A/J$, $J_{0}^{\perp}/J$, $\theta$, $\alpha$, $d/a$ in our spin model—one can generate diverse Meron Kekulé lattices characterized by different values of $l$, $l^{\prime}$, and $L$.
One straightforward avenue is to alter the twist angle ($\theta$), which leads to the expansion of $L$ proportionally to $L\approx\sqrt{3}a/\theta$ ($\theta$: in radian). This expansion enlarges local AA stacking structures, inducing corresponding expansions in AFM patches. We observe that $r$ expands at a rate commensurate with $L$ (i.e., $r\propto L$), as shown by the red line in Figure 4b. The $r$ expansion results in an increase in $l$, which is directly correlated with $r$, and subsequently $l^{\prime}$ through the relation $l^{\prime}=L-2l$. However, the expansion rates of $l$ and $l^{\prime}$ do not precisely overlap with that of $L$ (the blue and green lines in Figure 4b, respectively). Their linear behaviors are well-captured by the fitting lines: $r=0.28L+0.05$, $l=0.31L-1.04$, and $l^{\prime}=0.38L+2.09$, respectively. As a result, the expansion of $L$ results in the modulation of $\lambda$ (the inset in Figure 4b). It’s worth mentioning that the discrete alternation of $l$ and $l^{\prime}$ here originates from the discontinuous nature of the underlying honeycomb lattice, while the continuous behavior of $r$ is attributed to our determination of the AFM patch through the interpolation of the interlayer exchange energy map. For detailed methods for $r$ determination, refer to the Supplementary Information.
Alternatively, the adjustment of $r$ can be achieved by modifying the decay rate parameter of the interlayer exchange coupling ($\alpha$), as illustrated in Figure 4c. In this approach, $r$ can be continuously decreased by increasing $\alpha$ (the red line in Figure 4c). This leads to the corresponding contraction of $l$ with the alignment between $r$ and $l$ (the blue line in Figure 4c). Here, the decrease of $r$ through the increase of $\alpha$ is attributed to a significant suppression of $J_{ij}^{\perp}$ for a large value of $|\bm{r}_{ij}^{tb}|$ away from the center of an AFM patch. Similarly, reducing the ratio $d/a$ also reduces $r$ and contracts $l$ (for detailed analysis, refer to the Supplementary Information). It’s worth mentioning that within these two scenarios, $r$ and $l$ can be adjusted by keeping $L$ intact, in contrast to the former approach of adjusting $\theta$.
Instead of altering $r$, one can manipulate $l$ by adjusting the ratio between the strengths of the interlayer exchange interactions ($J_{0}^{\perp}$) and intralayer exchange interactions ($J$). As illustrated in Figure 4d, an increase in $J_{0}^{\perp}/J$ initially leads to an augmentation of $l$. However, as $J_{0}^{\perp}/J$ becomes significant, $l$ reaches its maximum value $l_{\textrm{max}}\approx r$. Further changes in $J_{0}^{\perp}/J$ do not affect $l$ thereafter. We attribute this intricate behavior to the interplay between two magnetic interactions influencing merons: (i) attractive Coulomb interactions between merons and antimerons with opposite vorticities 59, induced by the intralayer exchange interactions, and (ii) an effective moiré field generated by the bottom layer through interlayer exchange interactions 47. The former interaction tends to reduce $l$ to minimize the Coulomb energy of meron-antimeron pairs. In contrast, the latter interaction favors $l=r$ to minimize the field energy associated with the interlayer exchange energy. We posit that the interplay between these two mechanisms, quantified by the ratio of the two couplings $J_{0}^{\perp}/J$, determines the precise equilibrium positions of the meron cores at a specific $l$, leading to the observed modulation of $l$.
To elucidate the interplay between the Coulomb interactions and the moiré field, we utilize an energy function derived from an effective continuum model (for the derivation, refer to the Supplementary Information). This energy function is expressed as:
$E(l)=$ $\;C_{1}J\ln\Bigg{[}\sqrt{2l/3L}\frac{\big{(}1-(l/L)^{6}\big{)}^{2}\big{(}1-(2l/L)^{6}\big{)}}{\big{(}1+(\sqrt{3}l/L)^{6}\big{)}^{2}}\Bigg{]}$
$+\frac{C_{2}J_{0}^{\perp}}{a^{2}}(l-r)^{2}.$ (8)
Here, the first term describes the Coulomb interactions ranging up to the nearest-neighbor supercells, and the second term represents the moiré field. The constants $C_{1}$ and $C_{2}$ are numerical factors on the order of unity. To represent the relative strengths of the two interactions, we introduce a dimensionless ratio: $p=(2J_{0}^{\perp}/J)(C_{2}/C_{1})(L^{2}/a^{2})$, where a small $p$ value indicates that the Coulomb interactions dominate over the moiré field, and vice versa. We observe that when $p$ is small, $E(l)$ displays concave characteristics throughout all $l$ values (Figure 4e). In contrast, when $p$ is large, $E(l)$ exhibits a local minimum near $l\approx r$ (e.g., $l\approx 0.25L$ for $r=0.25L$ in Figure 4e), which stabilizes the MK state by positioning the meron cores at this location. In this regime, we determine the precise location of the meron cores by minimizing $E(l)$ with respect to $l$. Our analysis showcases that the resulting $l_{*}$ captures the qualitative features of the observed $l$ values (as depicted by the dashed lines in Figure 4d), suggesting the applicability of this effective model. For the detailed fitting method, refer to the Methods.
The effective potential $E(l)$ provides insights into the underlying mechanism supporting the stability of merons within the MK state. In conventional magnets, merons tend to be unstable due to Coulomb interactions, undergoing pair annihilation during energy relaxation processes. However, in our twisted magnets, the moiré field may mitigate this instability by confining the meron cores near the AFM domain boundary. Any deviation of the cores from this boundary results in an energy penalty, preventing pair annihilation. This protective effect of the moiré field is notably strengthened with increased interlayer exchange coupling or reduced twist angles, suggesting that the stabilization of the MK state occurs in conditions favoring strong intralayer exchange or small twist angles. Our numerical simulations validate this trend, aligning with predictions from the effective model (illustrated by the dashed line in Figure 4f). Consequently, we posit that adjusting the twist angle or interlayer exchange coupling strength can effectively stabilize the MK state.
@@
Based on our theoretical investigations, as depicted in Figure 4, we propose a methodology for realizing diverse Meron Kekulé lattice structures. Specified examples of such structures are illustrated in Table 1. Firstly, one can create Meron Kekulé lattices at various length scales ($l$, $l^{\prime}$, and $L$) by manipulating the size of the moiré superlattice ($L$) through adjustments in the twist angle ($\theta$) 60, 61, 62. Another method involves modifying the specifics of interlayer exchange coupling ($\alpha$, $d/a$, and $J_{0}^{\perp}/J$), which enables the manipulation of $l$ and $l^{\prime}$ while keeping $L$ constant. Experimentally, the latter approach can be implemented using different antiferromagnetic materials or external controls such as gate voltage, intercalation, adsorption, strain, or surface adsorption 63. The resulting Kekulé lattices display a wide range of distortion, as indicated by $\lambda$ values ranging from 0.05 to 0.34. This high degree of adaptability underscores a unique characteristic of the Meron Kekulé lattice, providing a promising avenue for experimental validation of its existence.
In real magnetic systems, the adaptability of the Meron Kekulé lattice might be limited by its stability conditions. For instance, our analysis reveals that insufficient single-ion anisotropy energy results in the decay of the MK state into the MD state through pair annihilation. Further details are available in the Supplementary Information. Additionally, it is crucial to consider material-specific properties that can impose constraints. For example, we assume that material values of $\alpha$ might range from 5 to 30, where $\alpha=5$ corresponds to $J_{ij}^{\perp}/J_{0}^{\perp}=0.47$ at the second-nearest neighbor distance between the top and bottom layers in the AA stacking, while $\alpha=30$ yields $J_{ij}^{\perp}/J_{0}^{\perp}=0.01$ at the same distance. This range aligns with previous ab-initio calculations for monolayer CrBr3 on substrate MnPSe3, estimating $\alpha\approx 10$ with $d/a\approx 1.89$ 51. Moreover, typical van der Waals magnetic materials exhibit $d/a\approx$ 1 to 2 and $J_{0}^{\perp}/J\approx$ 0.1 to 0.3 56. These material-specific values serve as constraining factors that potentially limit the adaptability of the Meron Kekulé lattice.
It is also worth noting that, unlike in our theoretical framework, the parameters that characterize the spin system may not be independently manipulable in real magnetic systems under experimental constraints. One plausible scenario is that reducing $d$ may lead to a simultaneous decrease in $d/a$ and an increase in $J_{0}^{\perp}/J$. In this context, our analysis indicates a potential nonmonotonic behavior of $l$; it may initially show a decreasing trend due to the high value of $J_{0}^{\perp}/J$, as illustrated in Figure 4d, followed by an increasing trend due to the enlargement of $r$, as suggested by Figure 4b. Predicting this behavior accurately presents an intriguing prospect for future research, requiring precise modeling of interlayer magnetic coupling through ab-initio calculations 51, 64, 65, 37, 54, 66. While this task is beyond the scope of the current study, we emphasize that the observed susceptibility of Kekulé structures holds promise for future experimental validation.
@@
where $\bm{B}=B\hat{\bm{z}}$ denotes external magnetic fields in the out-of-plane magnetization. The initial MD and MK states, as presented in Figures 3 and 4, are relaxed under external magnetic fields with continuously varying strength $B$. From the relaxed spin configurations for each state, the magnetic susceptibility, defined as $\chi\equiv dM/dB$, is calculated as a function of $B$, where $M$ is the average magnetization in the out-of-plane magnetization, defined as $M=\frac{1}{N}\sum_{i,l}\bm{S}_{i}^{l}\cdot\hat{\bm{z}}$. Figure 5ab presents the magnetic susceptibility $\chi$ of the MD state. Starting from the zero-field point $B=0$, $\chi$ initially decreases as $|B|$ increases, while it shows an upturn in $|B|>2.22$ meV, exhibiting shoulder-like features around $|B|=4.88$ meV (Figure 5a). Beyond these shoulders, $\chi$ sharply drops, and subsequently vanishes at $|B|>B_{\textrm{sat.}}=6.86$ meV (Figure 5b), where the magnetization saturates at $|M|=1$. We observe that despite this intricate behavior of $\chi$, the MD state persists throughout the entire field range, which is highlighted by the same yellow color.
@@
Figure 5cd presents the magnetic susceptibility $\chi$ of the MK state. In the low-field regime ($|B|<B_{c1}=0.52$ meV; colored by red), $\chi$ exhibits an increasing tendency with $|B|$ in contrast to the MD state. This behavior is attributed to the canted spin configurations in the core region that enhance the magnetization induced by the field. However, entering into the intermediate-field regime ($B_{c1}\leq|B|<B_{c2}=5.62$ meV; colored by blue), $\chi$ undergoes a sharp drop and subsequently exhibits a decreasing tendency similar to the MD state. Our findings reveal that at the threshold $|B|=B_{c1}$, the spin flipping occurs in the perpendicular direction for spins with out-of-plane components opposite to the field direction. The resulting state maintains in-plane swirling textures while lacking out-of-plane components in their Néel vectors. Consequently, this state retains only vorticities but loses half-skyrmion numbers characteristics, thus being named a “vortex state” (for the detailed spin configuration and skyrmion numbers analysis, refer to the Supplementary Information). Finally, upon entering the high-field regime ($|B|\geq B_{c2}$; colored by yellow), the vortex state transitions into the MD state through the annihilation of meron-antimeron pairs. At this regime, $\chi$ aligns with that of the MD state, as shown in the yellow region in Figure 5d. We highlight the two anomalous discontinuities at the transition points ($|B|=B_{c1}$ and $|B|=B_{c2}$), which are absent in the MD state, as discernible signatures of the formation of merons in the MK state. We propose experimental measurement of these anomalies to validate the Meron Kekulé lattice.
Despite our specific focus on the configuration depicted in Figure 3, the MK state can manifest in a diverse array of $2^{6}$ potential configurations, each defined by two possible vorticities ($p=\pm 1$) for the six meron cores within a moiré supercell. Among these configurations, only $2^{5}$ could exhibit variations in energy levels since global polarity reversal has no impact on the energy state. Our analysis indicates that despite the multitude of configurations, they maintain almost identical energy levels, deviating by less than $10^{-6}$ percentage from each other. Additionally, completely interchanging the positions of vortices and antivortices introduces a factor of $2$, while swapping the top and bottom layers introduces another factor of $2$, resulting in a total of $G=2^{8}$ potential variants for the MK state, all possessing the same energy level. This suggests that each supercell can assume one of the $G$ configurations with nearly identical energies. Consequently, a system comprising $N_{s}$ supercells can present a total of $G^{N_{s}}$ distinct configurations, underscoring the vast diversity and potential intricacy of the MK state. For a more detailed analysis, refer to the Supplementary Information.
The configurational diversity of the MK state suggests that it might be more favorable at finite temperatures compared to the ground MD state. To evaluate this, we calculate the expected occupation number of Meron hexads in each moiré supercell, denoted by $f_{\textrm{MH}}$, at temperature $T$. Considering two possibilities, the MD and MK configurations for each supercell, we express $f_{\textrm{MH}}$ as: $f_{\textrm{MH}}\simeq\frac{Ge^{-E_{\textrm{MK}}/k_{B}T}}{e^{-E_{\textrm{MD}}/k_{B}T}+Ge^{-E_{\textrm{MK}}/k_{B}T}}$, where $E_{\textrm{MK}}$ and $E_{\textrm{MD}}$ represent the energies of the MK and MD states from a single supercell, respectively. In this calculation, we assume the energy of each supercell is independent of the presence or absence of a Meron hexad (an approximation that might hold unless a significant number of Meron hexads occupy the supercells). By utilizing the calculated value, $E_{\textrm{MK}}-E_{\textrm{MD}}=21.44$ meV at $\theta=1.02$°, we determine $f_{\textrm{MH}}=0.32$ at $T=40$ K. Similarly, for $E_{\textrm{MK}}-E_{\textrm{MD}}=9.29$ meV at $\theta=0.50$°, estimated from the fitting, we derive $f_{\textrm{MH}}=0.49$ at $T=20$ K (for further details, refer to the Supplementary Information). These results imply that achieving the MK state at a finite temperature is feasible by either increasing the temperature or adjusting the twist angle at a small value 67. During our investigations, we also observed an alternative meron configuration characterized by either a single meron pair or double meron pairs in one of the two layers. In realistic settings, these configurations might appear in some of the supercells, acting as defects in the Meron Kekulé lattice. Investigating the effect of these defects presents an intriguing prospect for future research.
Discussion
We highlight that the stabilizing mechanism for the Meron Kekulé lattice is solely attributed to twist-induced lateral modulation of atomic stacking structures, irrespective of detailed characteristics in the interlayer exchange coupling. This contrasts with previous twist-engineering approaches for achieving topological spin textures, such as skyrmions 64, 65, 53, 68, 54, 55, 69, 70, 71 and merons 47, which relied on stacking-dependent interlayer magnetism confined to the CrX3 family (X = I, Cl, Br) 72, 65, 73, 74. Therefore, our approach presents an intriguing possibility for realizing topological spin textures in diverse vdW AFM materials 75, 56 independent of such material-specific properties.
@@ Discussion
We propose that our theory can be applied to CoPS3, which exhibits two essential factors for achieving the Meron Kekulé lattice: AFM intralayer exchange interactions and easy-plane magnetic anisotropy. Another potential application lies in XPS3 (X = Mn, Ni, Fe) and MnPSe3, where their intrinsic easy-axis magnetic anisotropy can be potentially altered into easy-plane anisotropy through experimental control, such as strain, gate voltage, or surface adsorption 76, 77, 78, 79, 80.
An intriguing research question to explore is how the formation of a Meron Kekulé lattice affects magnon band structures through its distorted lattice structure in vdW antiferromagnets. One potential outcome suggested by prior investigations is the creation of topological band structures by opening a band gap through hopping amplitude modulation 41, 42, 43. Furthermore, exploring phonon-like collective-excitation modes stemming from meron core vibrations within the Meron Kekulé lattice presents a compelling avenue for investigating topological band structures, similar to the approach used for lattice-forming skyrmion cores 81, 82.
@@
Methods
Relaxation of spin configurations.
To achieve relaxed spin configurations, we employed an “iterative optimization method” 54, 55, 47, 83. We initially generated random spin configurations represented by normalized vectors $\bm{n}_{i}=\bm{S}_{i}/|\bm{S}_{i}|$ for a classical spin vector $\bm{S}_{i}$ at each $i$-th site. An effective magnetic field $\bm{h}_{i}$, acting on each $\bm{S}_{i}$, was computed as $\bm{h}_{i}\equiv\frac{\partial H[\{\bm{n}_{i}\}]}{\partial\bm{n}_{i}}$. Subsequently, we iteratively updated $\bm{n}_{i}$ by subtracting $\bm{h}_{i}$ from $\bm{n}_{i}$, resulting in a spin configuration with reduced magnetic energy at each iteration. The update rule is defined as:
$\bm{n}_{i}^{(s+1)}=\frac{\bm{n}_{i}^{(s)}-c\bm{h}_{i}^{(s)}}{\left|\bm{n}_{i}^{(s)}-c\bm{h}_{i}^{(s)}\right|},$ (10)
where $(s)$ denotes the iteration step, and $c$ controls the update rate. The iteration proceeds until $\bm{h}_{i}^{(s)}$ becomes parallel to $\bm{n}_{i}^{(s)}$, indicating that further updates will not alter the spin configuration and ensuring minimal magnetic energy. To guarantee numerical convergence, we employed the condition $\frac{1}{N}\sum_{i}|\bm{h}_{i}^{\perp}|<10^{-10}$, where $\bm{h}_{i}^{\perp}=\bm{h}_{i}-\bm{n}_{i}(\bm{n}_{i}\cdot\bm{h}_{i})$. This convergence criterion aligns with the stability condition in the Landau-Lifshitz-Gilbert equation governing the magnetic dynamics. Throughout the iterative updates, we applied periodic boundary conditions to maintain consistency across the boundaries of the moiré unit cell.
Determination of the phase diagram.
To obtain the phase diagram presented in Figure 2d, we used random initial configurations denoted as $\{\bm{n}_{i}^{(0)}\}=\{(\frac{u_{i}}{n_{i}},\frac{v_{i}}{n_{i}},\frac{w_{i}}{n_{i}})\}$, where $u_{i}$, $v_{i}$, and $w_{i}$ are random numbers within the range of $[-1,1]$ and $n_{i}=\sqrt{u_{i}^{2}+v_{i}^{2}+w_{i}^{2}}$. These initial configurations were relaxed using an iterative optimization method to obtain the relaxed configuration $\{\bm{n}_{i}\}$. The ground state was identified as the lowest energy configuration among $\{\bm{n}_{i}\}$ for each parameter set ($\theta,J$). From the resulting ground states, the order parameters $\Psi_{l}$, defined as
$\Psi_{l}=\left|\frac{1}{N_{l}}\sum_{i}\bm{N}_{i}^{l}\right|$ (11)
were calculated for each layer $l=t,b$. Here, $N_{l}$ is the number of spins in each layer $l$. The AFM phase is characterized by $\Psi_{t}=\Psi_{b}=1$, while the MD phase signifies $\Psi_{t}<1$ and $\Psi_{b}<1$. $\Psi_{t}$ was presented in Figure 2d. It was found that the difference between $\Psi_{t}$ and $\Psi_{b}$ is less than $10^{-13}$.
Computation of the vorticity and skyrmion densities.
To compute the vorticity and skyrmion density maps of the MK state, depicted in Figure 3cd, we performed interpolation of the Néel vectors representing the spin configuration from the original honeycomb lattice to a square lattice consisting of $N_{\textrm{sq}}$ sites. Subsequently, the vorticity and skyrmion densities were calculated using these interpolated Néel vectors via Eqs. (5) and (6), respectively, with $N_{\textrm{sq}}=201^{2}$. This method yielded consistent values for vorticity and skyrmion numbers, as demonstrated by obtaining $V_{1}=0.9985$ and $Q_{1}=-0.4926$ for the meron $\textrm{M}_{1}$ when employing $N_{\textrm{sq}}=1,001^{2}$. These values closely align with the theoretical expectations of $V_{1}=1$ and $Q_{1}=-1/2$. The positions of the meron cores, as depicted in Figure 4bd, were identified based on the peak locations of these maps.
Determination of various Kekulé structures from different parameter sets.
To determine $l$ of the MK state from different parameter sets, as presented in Figure 4bd, we utilized various spin configurations characterized by different $l$ values. These initial trial configurations were generated through the interpolation of the original MK state to attain diverse initial $l/L$ values ranging from 0.2 to 1.2 with an interval of 0.02. Subsequently, for each parameter set, the initial configurations were relaxed through the iterative optimization method to obtain stable spin configurations that potentially possess different $l$. From the relaxed spin configurations, an energy profile of the MK state, $E_{\textrm{simul.}}(l)$, was computed as a function of $l$. $E_{\textrm{simul.}}(l)$ typically exhibits a convex behavior with a local minimum. The predominant MK states for each parameter set were selected based on the minimized energy configuration, and the $l$ values were computed from these states. Similarly, the stabilization condition of the MK state, as presented in Figure 4f, was determined from the energy profile; the presence (absence) of the local minimum in this profile indicates the stable (unstable) nature of the MK state in each phase space. Further details are available in the Supporting Information.
Predictions made from the effective energy function.
To derive the fitting curve in Figure 4d, we identified a local minimum $l_{*}$ of $E(l)$ in Eq. (8) by solving the equation $\frac{dE(l_{*})}{dl}=0$ under the constraint $\frac{d^{2}E(l_{*})}{dl^{2}}>0$. In the calculations, the parameters $J=1$, $L=24.60$, and $a=0.4$ were used, and $r=7.22$ were adopted from the simulation. The fitting parameter $C\equiv C_{2}/C_{1}=19.74$ was employed. The fitting curve in Figure 4f was derived from the formula $J_{0}^{\perp}=D/L^{2}$ where $D=\frac{pJa^{2}C}{2}$. In the fitting, the parameters $J=1$, $a=0.4$, and $p=23.6$ were used, where the value of $p$ was determined by the threshold value of $p$ for the stabilization of the MK state as determined from our numerical investigations (further details are available in the Supporting Information). The fitting parameter $C$ was estimated as $C=21.33$, which is consistent with the estimation of $C=19.74$ from Figure 4d.

### v2

@@ Abstract
The cores of the merons and antimerons form six vertices of a hexagon along the boundary of the AFM patch (Figure 3a). We refer to this distinctive configuration as a “Meron hexad.” Despite its intricate texture, the Meron hexad attains an AFM domain array configuration that minimizes interlayer exchange energy across the bulk region (Figure 3e). Furthermore, the constituent meron cores of Meron hexads form a honeycomb-lattice-like pattern, where the distance between two nearest-neighbor meron cores from the same hexad is shorter than that between cores from adjacent Meron hexads (Figure 3a). This distorted structure is referred to as a Kekulé lattice with a Kekulé-O pattern 9. Thus, we term this distorted meron lattice a “Meron Kekulé lattice (MK).” The Kekulé structure of the MK state is characterized by two key length parameters: the intracell bond length within the same Meron hexad ($l$) and the intercell bond length across different Meron hexads ($l^{\prime}=L-2l$). The distinction between $l$ and $l^{\prime}$ indicates the deviation from the conventional honeycomb lattice, giving rise to an alternative Kekulé structure. To quantify this deviation, we introduce a dimensionless constant $\lambda$, referred to as the “Kekulé constant:”
$\lambda=\frac{l^{\prime}-l}{L}=1-\frac{3l}{L},$ (7)
where $\lambda=0$ and $\lambda>0$ indicate regular honeycomb and Kekulé lattices, respectively.
We illustrate how variations in the parameters of the spin model influence the values of $l$ and $l^{\prime}$, as well as the AFM patch size ($r$). Figure 4b shows the evolution of $l$ and $l^{\prime}$ as functions of $L$. We first observe a linear increase in the AFM patch size ($r$) with $L$, which is evident since the increase in $L$ signifies expansions in all local stacking regions, including AFM patches. The increase in $r$ correlates with an increase in $l$, as the minimization of interlayer exchange energy dictates $l=r$ 28. In a large $L$ regime, this relationship holds precisely; however, it progressively deviates in a small $L$ regime. This deviation from $l=r$ contributes to the enhancement of Kekulé distortion, represented by $\lambda$ (see the inset). The linear increase in $l^{\prime}$ directly follows the increases in $L$ and $l$, as described by the relationship $l^{\prime}=L-2l$.
We find that an increase in $\alpha$ leads to the contraction of $r$, resulting in a decrease in $l$ while adhering to the relationship $l\approx r$. Meanwhile, $l^{\prime}$ increases according to $l^{\prime}=L-2l$ with a fixed $L$ (Figure 4c). The contraction of $r$ is attributed to significant suppression of $J_{ij}^{\perp}$ for large values of $|\bm{r}_{ij}^{tb}|$ away from the center of an AFM patch (see the insets). Moreover, an increase in $J_{0}^{\perp}/J$ initially leads to an increase in $l$ when $J_{0}^{\perp}/J$ is sufficiently small (Figure 4d). However, $l$ reaches its maximum value $l_{\text{max}}\approx r$, and further changes in $J_{0}^{\perp}/J$ do not affect $l$. We also find that reducing the ratio $d/a$ lowers $l$ by decreasing $r$, similar to the effects of $\alpha$. Furthermore, varying $A/J$ does not alter $l$, while extreme values lead to a transition to other magnetic states. Further details regarding this analysis of $d/a$ and $A/J$ can be found in SI 52.
@@
Table 1 presents Meron Kekulé lattices at various length scales ($l$, $l^{\prime}$, and $L$). These structures can be created by manipulating the twist angle ($\theta$) 55, 56, 57 or by modifying the parameters of interlayer exchange coupling ($\alpha$, $d/a$, and $J_{0}^{\perp}/J$). The latter approach can be implemented using different antiferromagnetic materials or external controls such as gate voltage, intercalation, adsorption, strain, or surface modification 58. The resulting Kekulé lattices exhibit a wide range of distortion, with $\lambda$ values ranging from 0.05 to 0.34, highlighting the adaptability of the Meron Kekulé lattice.
We attribute the intricate behavior of $l$ observed in Figure 4bd to the interplay between two magnetic interactions influencing merons: (i) attractive Coulomb interactions between merons and antimerons with opposite vorticities 54, induced by the intralayer exchange interactions, and (ii) an effective moiré field generated by the bottom layer through interlayer exchange interactions 28. To elucidate their interplay, we utilize an energy function derived from an effective continuum model:
$E(l)=$ $\;C_{1}J\ln\Bigg{[}\sqrt{\frac{2l}{3L}}\frac{\big{(}L^{6}-l^{6}\big{)}^{2}\big{(}L^{6}-64l^{6}\big{)}}{L^{6}\big{(}L^{6}+27l^{6}\big{)}^{2}}\Bigg{]}$ (8)
$+\frac{C_{2}J_{0}^{\perp}}{a^{2}}(l-r)^{2}.$
Here, the first term describes the Coulomb interactions ranging up to the nearest-neighbor supercells, and the second term represents the moiré field. The constants $C_{1}$ and $C_{2}$ are numerical factors on the order of unity. The relative strength of these two interactions is represented by the dimensionless ratio:
$p=\frac{2C_{2}J_{0}^{\perp}L^{2}}{C_{1}Ja^{2}}.$ (9)
The value of $l$, where the MK state is stabilized, denoted as $l_{*}$, can be determined by minimizing $E(l)$ with respect to $l$. When $p$ is large, $l_{*}\approx r$ (e.g., $l\approx 0.25L$ for $r=0.25L$ in Figure 4e), indicating that the moiré field term dominates. Conversely, a reduction in $p$ enhances the role of the Coulomb interaction term, leading to a decrease in $l_{*}$, as shown by the shifts in the local minimum of $E(l)$ in Figure 4e. We assert that this competition plays a crucial role in the intricate behavior of $l$ observed in Figures 4bd. In particular, our analysis demonstrates that the $l_{*}$ value derived from $E(l)$ captures the qualitative features of the observed $l$ values in the simulation, supporting this assertion (the dashed lines in Figure 4d). For the derivation of $E(l)$ and the fitting method for $l$, refer to SI 52.
The effective potential $E(l)$ offers insights into the underlying mechanisms that support the stability of merons within the MK state. In conventional magnets, merons tend to be unstable due to attractive Coulomb interactions, which lead to pair annihilation during energy relaxation processes. However, in our twisted magnets, the moiré field mitigates this instability by imposing an energy penalty for the deviation of the meron cores from the AFM domain boundary, leading to the energetic stability of the MK state, shown by the convexity of $E(l)$ (Figure 4e). Furthermore, the effective model predicts that the stabilization of the MK state can be achieved by increasing either the interlayer exchange coupling ($J_{0}^{\perp}$) or the superlattice size ($L$), which is confirmed in our numerical simulations (Figure 4f).
@@
where $\bm{B}=B\hat{\bm{z}}$ denotes external magnetic fields in the out-of-plane magnetization. The initial MD and MK states, presented in Figures 3 and 4, are relaxed under external magnetic fields with continuously varying strength $B$. From the relaxed spin configurations for each state, the magnetic susceptibility, $\chi\equiv dM/dB$, is calculated as a function of $B$, where $M=\frac{1}{N}\sum_{i,l}\bm{S}_{i}^{l}\cdot\hat{\bm{z}}$ is the average magnetization in the out-of-plane magnetization. Figure 5ab presents $\chi$ of the MD state. Near $B=0$, $\chi$ initially decreases as $|B|$ increases, while it shows an upturn in $|B|>2.22$ meV, exhibiting shoulder-like features around $|B|=4.88$ meV (Figure 5a). Beyond these shoulders, $\chi$ sharply drops, and subsequently vanishes at $|B|>B_{\textrm{sat.}}=6.86$ meV (Figure 5b), where $M$ saturates to $|M|=1$.
@@
Figure 5cd presents $\chi$ of the MK state. In the low-field regime ($|B|<B_{c1}=0.52$ meV), $\chi$ exhibits an increasing tendency with $|B|$ in contrast to the MD state. This behavior is attributed to the canted spin configurations in the core region that enhance the magnetization induced by the field. However, entering into the intermediate-field regime ($B_{c1}\leq|B|<B_{c2}=5.62$ meV), $\chi$ undergoes a sharp drop and subsequently exhibits a decreasing tendency. Our findings reveal that at the threshold $|B|=B_{c1}$, the spin flipping occurs in the perpendicular direction for spins with out-of-plane components opposite to the field direction. The resulting state maintains in-plane swirling textures while lacking out-of-plane components in their Néel vectors. Consequently, this state retains only vorticities but loses half-skyrmion numbers characteristics, thus being named a “vortex state” (for the detailed analysis of this state, refer to SI 52). Finally, upon entering the high-field regime ($|B|\geq B_{c2}$), the vortex state transitions into the MD state through the annihilation of meron-antimeron pairs. At this regime, $\chi$ aligns with that of the MD state. We highlight the two anomalous discontinuities at the transition points ($|B|=B_{c1}$ and $|B|=B_{c2}$) as discernible signatures of the formation of merons in the MK state.
Aside from the configuration in Figure 3, the MK state can manifest in a diverse array of $2^{6}$ potential configurations, each defined by two possible polarities ($p=\pm 1$) for the six meron cores within a Meron hexad. This configurational diversity suggests that the MK state may be more favorable at finite temperatures compared to the MD state. We find that the expected occupation number of Meron hexads in each supercell, denoted as $f_{\text{MH}}$, is $f_{\text{MH}}=0.32$ at $\theta=1.02^{\circ}$ and $T=40$ K, while $f_{\text{MH}}=0.49$ at $\theta=0.50^{\circ}$ and $T=20$ K. These results imply that achieving the MK state can be facilitated by either increasing the temperature or adjusting the twist angle to a lower value 59. Further details can be found in SI 52.
@@
We propose that our theory can be applied to CoPS3, which exhibits two essential factors for achieving the Meron Kekulé lattice: AFM intralayer exchange interactions and easy-plane magnetic anisotropy. Another potential application lies in XPS3 (X = Mn, Ni, Fe) and MnPSe3, where their intrinsic easy-axis magnetic anisotropy can be potentially altered into easy-plane anisotropy through experimental control, such as strain, gate voltage, or surface adsorption 60, 61, 62, 63, 64.
@@
An intriguing research question to explore is how the formation of a Meron Kekulé lattice affects magnon band structures through its distorted lattice structure in vdW antiferromagnets. One potential outcome suggested by prior investigations is the creation of topological band structures by opening a band gap through hopping amplitude modulation 22, 23, 24. Furthermore, exploring phonon-like collective-excitation modes stemming from meron core vibrations within the Meron Kekulé lattice presents a compelling avenue for investigating topological band structures, similar to the approach used for lattice-forming skyrmion cores 65, 66.